OpenStudy (anonymous):

In need of help :/ ..question > If you sleep 7 hours every night how many hours are you AWAKE in a year? Tips: There are 24 hours in a day. There are 365 days in a year

OpenStudy (anonymous):

24 - 7 = 17 17 * 365 = ? http://lmgtfy.com/?q=17+*+365

OpenStudy (anonymous):

Thank you, i was confused how to set it up :)

OpenStudy (anonymous):

24*365 =8760 7*365=2555 8760-2555=6205 You are awake for 6205 hours a year :)